A bag contains 10 blue, 6 red, and 4 green rubber balls. If you select a ball at random from the bag, what is the probability of not choosing a green ball?

A. 1/5
B. 4/5
C. 5/6
D. 9/10
Why can you eliminate choice D in this question?

Found 2 solutions by Edwin McCravy, Fombitz:
Answer by Edwin McCravy(20086)


A bag contains 10 blue, 6 red, and 4 green rubber balls. If you select a ball
at random from the bag, what is the probability of not choosing a green ball? 

A. 1/5 
B. 4/5
C. 5/6
D. 9/10 

To not choose a green ball means to select a non-green ball.  There are
16 non-green balls out of the 20. So the probability is 16/20 which reduces
to 4/5, so the correct answer is B.

----------------------------------------------------------

Why can you eliminate choice D in this question?

Because that's what the answer would be if there had been 10 blue, 8 red,
and 2 green rubber balls. 

Edwin

Answer by Fombitz(32388)

The total number of balls available is 10+6+4=20 balls.
The probability of choosing a blue ball is 10/20.
The probability of choosing a red ball is 6/20.
The probability of choosing a green ball is 4/20.
P(green or not green) = 1, since the ball I choose must be either green or not green.
P(green) + P(not green) = P(green or not green)
4/20 + P(not green) = 1
P(not green) = 16/20 = 4/5
The answer is B.
You can eliminate D because none of the probability can sum to 18 out of 20(9/10). The only possible choices are 16/20, 14/20, or 10/20.
You can also eliminate C because it's not possible to sum to repeating fraction from the possible probabilities (0.5,0.3,0.2).